OpenStudy (elisaneedshelp):

Please and thank you :( (I have a deadline today in 3 hours) A Web music store offers two versions of a popular song. The size of the standard version is 2.3 megabytes (MB). The size of the high-quality version is 4.7 MB. Yesterday, there were 1120 downloads of the song, for a total download size of 3128 MB. How many downloads of the standard version were there?

OpenStudy (anonymous):

?

OpenStudy (elisaneedshelp):

?? ...

OpenStudy (elisaneedshelp):

@jigglypuff314

jigglypuff314 (jigglypuff314):

The total number of standard versions plus high quality versions is 1120 downloads right? :) do you think you can make that into an equation?

OpenStudy (elisaneedshelp):

I put 2.3+(1120)-2.3x+4.7x=3128 I took a problem i tried to do before as an example.

OpenStudy (elisaneedshelp):

but i dont know

jigglypuff314 (jigglypuff314):

hmm you should try the system of equations s + h = 1120 2.3s + 4.7h = 3128

OpenStudy (elisaneedshelp):

Oh, they didnt explain it to us that way~! ._.

jigglypuff314 (jigglypuff314):

do you know how to do that kind of system of equations? :)

OpenStudy (elisaneedshelp):

yeah (:

OpenStudy (elisaneedshelp):

I got 890 and it was correct. Ill do the rest of them by myself. Can i tag you in another question after? or do you have class

jigglypuff314 (jigglypuff314):

aye I have a major project due in 8 hours ... that Is didn't start

OpenStudy (elisaneedshelp):

lol oh :c and how do i do system of equations when there are only 3 numbers in the problem i have to do now?

jigglypuff314 (jigglypuff314):

s + h = 1120 2.3s + 4.7h = 3128 s = 1120 - h 2.3(1120 - h) + 4.7h = 3128 distribute and solve for h then plug back into s = 1120 - h to get s